OpenStudy (anonymous):

Factoring x²=-2x+33 ?

OpenStudy (anonymous):

just solve x^2+2x-33=0 and get two roots x1, x2

OpenStudy (anonymous):

help please :/

OpenStudy (anonymous):

x²=-2x+33 x^2+2x=33 x(x+2)=33 x=33 or x+2 =33

OpenStudy (dumbcow):

set it equal to zero x^2 +2x -33 = 0 for this to factor, 2 factors of -33 must add up to 2 none exist, this is not factor-able to solve for x, use quadratic formula x = -b +- sqrt(b^2 -4ac) / 2a

OpenStudy (anonymous):

and then what ?

OpenStudy (anonymous):

(x-x1)(x-x2)=0